Dawson Toys, Ltd., produces a toy called the Maze. The company has recently established a standard cost system to help control costs and has established the following standards for the Maze toy:

Direct materials: 8 microns per toy at $0.30 per micron
Direct labor: 1.3 hours per toy at $6.70 per hour
During July, the company produced 5,000 Maze toys. Production data for the month on the toy follow:
Direct materials: 72,000 microns were purchased at a cost of $0.29 per micron. 22,000 of these microns were still in inventory at the end of the month.
Direct labor: 6,900 direct labor-hours were worked at a cost of $49,680.

Compute the following variances for July: (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations. Round final answer to the nearest whole dollar.)

a. The materials price and quantity variances.
b. The labor rate and efficiency variances.
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (0.30-0.29) *72,000

= 720 F

Materials Quantity Variance = (SQ-AQ) *SP

= (5000*8 - 50,000) * 0.30

= 3000 U

Labor Rate Variance = (SR-AR) * AH

= (6.70 - 49,680/6900) * 6900

= 3450 U

Labor efficiency variance = (SH-AH) * SR

= (5000*1.3 - 6900) * 6.70

= 2680 U